# Failure of the measure contraction property via quotients in higher-step sub-Riemannian structures

### v1

@@ Abstract.
We investigate the validity of the synthetic Ricci curvature lower bound known as the measure contraction property ($\mathrm{MCP}$) for sub-Riemannian structures beyond step two. We show that whenever the distance function is not Lipschitz in charts, the $\mathrm{MCP}$ may fail. This occurs already in fundamental examples such as the Martinet and Engel structures.
Central to our analysis are new results on the stability of the local $\mathrm{MCP}$ under quotients by isometric group actions for general metric measure spaces, developed under a weaker variant of the essential non-branching condition which, in contrast with the classical one, is implied by the minimizing Sard property in sub-Riemannian geometry.
Since the $\mathrm{MCP}$ is preserved under blow-ups, we focus on Carnot homogeneous spaces, proving that $\mathrm{MCP}$ descends to suitable quotients. As a byproduct, any structure whose tangent at some point admits a quotient to Martinet fails the $\mathrm{MCP}$. We also obtain a computation-free proof that the Grushin plane shares the Heisenberg group’s $\mathrm{MCP}$. Applications include a detailed analysis of validity and failure of the $\mathrm{MCP}$ for Carnot groups of low dimension.
Our results suggest a conjecture on the failure of the $\mathrm{MCP}$ in presence of Goh abnormal geodesics satisfying the strong generalized Legendre condition.
@@ 1.1. Failure of the measure contraction property for Martinet
The proof of Theorem 1.1 relies upon the optimal synthesis of the Martinet structure, which was obtained in [ABCK97]. We note that we work with the Ohta’s $\mathrm{MCP}$, which is the weakest variant of the $\mathrm{MCP}$, so that any other formulation fails as well. Theorem 1.1 is actually the last result we prove in this paper, but it is the starting point and motivation for the rest of our work.
@@ 1.2. Measure contractions for quotients
However, as anticipated, there are non-trivial technical challenges preventing the application of Theorem 1.2 to quotients such as (1).
@@ 1.3. Dealing with Item i: $\delta$-essentially non-branching
Let $(X,\mathsf{d},\mathsf{m})$ be a $\delta$-essentially non-branching metric measure space satisfying the $\mathrm{MCP}_{\mathrm{loc}}(K,N)$ for some $K\in\mathbb{R}$ and $N\in[1,\infty)$. Let $G$ be a discrete group of metric measure isometries of $(X,\mathsf{d},\mathsf{m})$ acting properly and freely. Then, the metric measure quotient $(X^{*},\mathsf{d}^{*},\mathsf{m}^{*})$ satisfies the $\mathrm{MCP}_{\mathrm{loc}}(K,N)$. Furthermore, if $(X,\mathsf{d},\mathsf{m})$ is $\delta$-essentially non-branching, then $(X^{*},\mathsf{d}^{*},\mathsf{m}^{*})$ is $\delta$-essentially non-branching.
@@ 1.4. Dealing with Item ii: local measure contractions for quotients
Let $(X,\mathsf{d},\mathsf{m})$ be a $\delta$-essentially non-branching metric measure space satisfying the $\mathrm{MCP}_{\mathrm{loc}}(K,N)$ for some $K\in\mathbb{R}$ and $N\in[1,\infty)$. Let $G$ be a connected, compact, abelian Lie group acting freely by metric measure isometries on $X$. Then, the metric measure quotient $(X^{*},\mathsf{d}^{*},\mathsf{m}^{*})$ satisfies the $\mathrm{MCP}_{\mathrm{loc}}(K,N)$.
@@ 1.5. Dealing with Item iii: the factorization argument
We remark that while all results presented so far hold for general metric measure spaces, what follows is specific for quotients between Carnot homogeneous spaces. In this case, we can exploit the stratified nature of these spaces to replace the naive two-step factorization outlined in Item ii with a more elaborate construction. Namely, we factor (1) via multiple pairs of local isometries and quotients by compact abelian groups such that, at each factor, both Theorems 1.4, LABEL: and 1.5 can be applied. This is the content of the next result (corresponding to Theorem 6.16, to which we refer to for additional comments and remarks).
@@ 1.6. Measure contraction properties beyond step 2
Suppose that a Carnot homogeneous space $\mathbb{M}=\mathbb{G}/\mathbb{H}$ admits a quotient to another Carnot homogeneous space $\tilde{\mathbb{M}}=\mathbb{G}/\tilde{\mathbb{H}}$. Assume that the minimizing Sard property holds for all factors $\mathbb{M},\mathbb{M}_{1},\dots,\mathbb{M}_{s-1}$ of the map $\pi:\mathbb{M}\to\tilde{\mathbb{M}}$ in Theorem 6.16 (this is the case, for example, if the minimizing Sard property holds for all Carnot homogeneous spaces of step $\leq s$, where $s$ is the step of $\mathbb{M}$). Then, if $\mathbb{M}$ satisfies the $\mathrm{MCP}(K,N)$ for some $K\in\mathbb{R}$ and $N\in[1,\infty)$, the space $\tilde{\mathbb{M}}$ also satisfies the $\mathrm{MCP}(K,N)$.
@@
It is not trivial to determine whether a Carnot group admits a quotient to the Martinet structure. This is a purely algebraic condition, independent on the metric geometry of the space. Quite unexpectedly, we prove in the next result that admitting a quotient to Martinet structure is equivalent to admit a quotient to the Engel group (it corresponds to Theorem 6.13).
@@
A characterization of Carnot groups admitting a quotient to the Engel group appeared in [Mar14, Prop. 5.1], in terms of a special basis of the Lie algebra. In contrast, condition (7) in Theorem 1.11 is intrinsic, and it is key for the next applications. See Remark 6.14 for a sufficient condition in terms of the dimension of $\mathbb{G}$.
@@ 1.7. Applications to Carnot groups in low dimensions
We apply our results to Carnot groups in low dimension. Recall that a Carnot group is indecomposable if its Lie algebra is not the direct sum of two non-trivial Lie algebras. A classification of indecomposable Carnot groups of dimension $\leq 7$ is provided in [LT22, Gon98], to which we refer for details.
Goh-Legendre geodesics are a class of abnormal geodesics satisfying second-order necessary conditions for minimality, see Definition A.1. On the one hand, if we are able to rule out the existence of Goh-Legendre geodesics on a given Carnot group, we can pair the results in [AL09] and [BR20a] to prove the validity of the $\mathrm{MCP}$ for that group, see Section 7.2 for details.
@@
Remark 1.13.
The smallest $N$ for which the conclusion of Corollary 1.12 holds true is called the curvature exponent of the Carnot group, see [Rif13]. It is known that this number must be greater than or equal to the so-called upper geodesic dimension, which in turn is strictly greater than the Hausdorff dimension of the Carnot group, see [Zha25, GZ24, Riz16].
@@
Among the indecomposable Carnot groups of dimension $\leq 7$, listed in Table 1:
@@ 5 $N_{5,3,1}$, $N_{5,3,2}$ $N_{5,2,3}$ $N_{5,2,1}$
Table 1. Each entry is an indecomposable Carnot group of dimension $\leq 7$, see [LT22, Gon98]. Red: admits a quotient to the Martinet structure. Green: does not have non-trivial Goh-Legendre geodesics. Black: does not admit a quotient to Martinet and admits non-trivial Goh-Legendre geodesics (see Remark 1.15). : conditional on the validity of the minimizing Sard property (see Remark 1.16). : admits non-trivial Goh geodesics but no Goh-Legendre ones (see Remark 7.7).
@@
We do not know whether the $\mathrm{MCP}$ is satisfied by the remaining Carnot groups in Table 1. They do not admit a quotient to the Martinet structure, so Theorem 1.9 cannot be applied. They do have non-trivial Goh–Legendre geodesics, so we cannot use Corollary 1.12 either. Among these groups, it seems that $247H$, $247H_{1}$, $247K$, $247N$, and $247P_{1}$ admit a quotient to $N_{6,2,6}$ or $N_{6,3,1a}$. If one could prove by other means that the $\mathrm{MCP}$ fails for these two groups, it would imply the failure of the $\mathrm{MCP}$ for all the others.
Remark 1.16 (Conditionality on Sard).
@@
We end with a straightforward result for decomposable groups (see Corollary 7.8).
@@ 1.8. Closing thoughts
If a sub-Riemannian metric measure space $(M,\mathsf{d},\mathsf{m})$ does not have non-trivial Goh-Legendre geodesics, then, by the results of Agrachev and Lee in [AL09], the distance function is locally Lipschitz in charts out of the diagonal. Consequently, thanks to the work of Badreddine and Rifford in [BR20a], and assuming that the structure is real-analytic, it must, locally, satisfy the $\mathrm{MCP}$. If non-trivial Goh-Legendre geodesics are present, the $\mathrm{MCP}$ (and local Lipschitz regularity) may fail, and indeed they do fail for many examples in Table 1.
@@
Let $(M,\mathsf{d},\mathsf{m})$ be a sub-Riemannian metric measure space. If there is a non-trivial strong-Goh-Legendre geodesic $\gamma$, then $\mathsf{d}:M\times M\to\mathbb{R}_{+}$ fails to be Lipschitz in charts in a neighbourhood of $\gamma$, and the $\mathrm{MCP}(K,N)$ does not hold for all $K\in\mathbb{R}$ and $N\in[1,\infty)$.
@@
Using Corollary 1.17, we can produce Carnot groups that have non-trivial Goh–Legendre geodesics and yet satisfy the $\mathrm{MCP}$, by taking products of factors with no non-trivial Goh–Legendre geodesics. Such geodesics are not strong Goh–Legendre ones though, supporting the conjecture.
The absence of Goh-Legendre geodesics also implies, for real-analytic structures, that the corresponding distance function is subanalytic, as proven by Agrachev and Gauthier in [AG01]. On the contrary, akin to what happens for the $\mathrm{MCP}$ and, of course, local Lipschitzianity, subanalyticity is lost when a suitable Martinet quotient do exist. All these properties – local Lipschitzianity, $\mathrm{MCP}$ and subanalyticity – (or their failure) seem to be a manifestation of the compactness of the set of normal geodesics of a fixed length (or lack thereof). However, a direct connection between them is currently missing.
@@ 1.9. Structure of the paper
Section 2 contains an overview of optimal transport on metric measure spaces and their quotients by isometric group actions. In Section 3, we introduce the weaker essentially non-branching condition that we require, and discuss its relation with the $\mathrm{MCP}$. We then address the stability of the $\mathrm{MCP}$ under discrete and compact group actions in Sections 4 and 5, respectively. In Section 6, we present Carnot homogeneous spaces and we prove the factorization argument. We put everything together in Section 7, proving our main theorem concerning the $\mathrm{MCP}$ of quotients of Carnot homogeneous spaces and its consequences. We end with Section 8, where we prove the failure of the $\mathrm{MCP}$ for the Martinet structure.

### v2

@@ Abstract.
We prove that the synthetic Ricci curvature lower bound known as the measure contraction property ($\mathrm{MCP}$) can fail in sub-Riemannian geometry. This may happen beyond step two, if the distance function is not Lipschitz in charts, and it already occurs in fundamental examples such as the Martinet and Engel structures.
Central to our analysis are new results, of independent interest, on the stability of the local $\mathrm{MCP}$ under quotients by isometric group actions for general metric measure spaces, developed under a weaker variant of the essential non-branching condition which, in contrast with the classical one, is implied by the minimizing Sard property in sub-Riemannian geometry.
As an application, we find sub-Riemannian structures with pre-medium-fat distribution that do not satisfy the $\mathrm{MCP}$, answering a question raised in [Rif23]. Finally, and quite unexpectedly, we show that ideal sub-Riemannian structures can fail the $\mathrm{MCP}$, and this actually happens generically for rank greater than $3$ and high dimension.
@@ 1.1. Failure of the measure contraction property for Martinet
The proof of Theorem 1.1 relies upon the optimal synthesis of the Martinet structure, which was obtained in [ABCK97]. Theorem 1.1 is actually the last result we prove in this paper, but it is the starting point and motivation for the rest of our work.
@@ 1.2. Measure contractions for quotients
However there are non-trivial challenges preventing the application of Theorem 1.2 to (1).
@@ 1.3. Dealing with Item i: $\delta$-essentially non-branching
Let $(X,\mathsf{d},\mathsf{m})$ be a metric measure space satisfying the $\mathrm{MCP}_{\mathrm{loc}}(K,N)$ for some $K\in\mathbb{R}$ and $N\in[1,\infty)$. Let $G$ be a discrete group of metric measure isometries, acting properly and freely. Then, the metric measure quotient $(X^{*},\mathsf{d}^{*},\mathsf{m}^{*})$ satisfies the $\mathrm{MCP}_{\mathrm{loc}}(K,N)$. If $(X,\mathsf{d},\mathsf{m})$ is $\delta$-essentially non-branching, then $(X^{*},\mathsf{d}^{*},\mathsf{m}^{*})$ is $\delta$-essentially non-branching.
@@ 1.4. Dealing with Item ii: local measure contractions for quotients
Let $(X,\mathsf{d},\mathsf{m})$ be a $\delta$-essentially non-branching metric measure space satisfying the $\mathrm{MCP}_{\mathrm{loc}}(K,N)$ for some $K\in\mathbb{R}$ and $N\in[1,\infty)$. Let $G$ be a connected, compact, abelian Lie group of metric measure isometries, acting freely. Then, the metric measure quotient $(X^{*},\mathsf{d}^{*},\mathsf{m}^{*})$ satisfies the $\mathrm{MCP}_{\mathrm{loc}}(K,N)$.
@@ 1.5. Dealing with Item iii: the factorization argument
While, all results presented so far hold for general metric measure spaces, what follows is specific for quotients between Carnot homogeneous spaces. In this case, we can exploit their stratified nature to replace the naive two-step factorization outlined in Item ii with a more elaborate construction. Namely, we factor (1) via multiple pairs of local isometries and quotients by compact abelian groups such that, at each factor, both Theorems 1.4, LABEL: and 1.5 can be applied. This is the content of the next result (corresponding to Theorem 6.16, to which we refer to for additional comments and remarks).
@@ 1.6. Measure contraction properties beyond step 2
Suppose that a Carnot homogeneous space $\mathbb{M}=\mathbb{G}/\mathbb{H}$ admits a quotient to a Carnot homogeneous space $\tilde{\mathbb{M}}=\mathbb{G}/\tilde{\mathbb{H}}$. Assume that the minimizing Sard property holds for all factors $\mathbb{M},\mathbb{M}_{1},\dots,\mathbb{M}_{s-1}$ of the map $\pi:\mathbb{M}\to\tilde{\mathbb{M}}$ in Theorem 6.16 (this is the case, for example, if the minimizing Sard property holds for all Carnot homogeneous spaces of step $\leq s$, where $s$ is the step of $\mathbb{M}$). Then, if $\mathbb{M}$ satisfies the $\mathrm{MCP}(K,N)$ for some $K\in\mathbb{R}$ and $N\in[1,\infty)$, the space $\tilde{\mathbb{M}}$ also satisfies the $\mathrm{MCP}(K,N)$.
@@
It is not trivial to determine whether a Carnot group admits a quotient to the Martinet structure. Quite unexpectedly, we prove in the next result that admitting a quotient to Martinet structure is equivalent to admit a quotient to the Engel group, and that both conditions are purely algebraic depending only on the underlying stratified group. This corresponds to Theorem 6.13.
@@
A characterization also appeared in [Mar14, Prop. 5.1], in terms of a special basis of the Lie algebra. In contrast, condition (7) in Theorem 1.11 is intrinsic, and it is key for the next applications. See Remark 6.14 for a sufficient condition in terms of the dimension of $\mathbb{G}$.
Different choices of scalar product on the first stratum of a stratified group may correspond to non-isometric Carnot groups (see Section 6). Theorem 1.11 and Theorem 1.9 yield the failure of the $\mathrm{MCP}$ for any Carnot structure on the same stratified group satisfying one of the equivalent conditions of Theorem 1.11. Such structures are all bi-Lipschitz equivalent, but the $\mathrm{MCP}$ is not preserved under bi-Lipschitz equivalence. For example, the $\ell^{p}$ Heisenberg group for $p>2$ does not satisfy any $\mathrm{MCP}$ [BT23]. We are not able to produce sub-Riemannian Carnot structures on the same stratified group, one satisfying the $\mathrm{MCP}$ and the other failing it.
@@ 1.7. Applications to Carnot groups in low dimensions
Recall that a stratified group is indecomposable if its Lie algebra is not the direct sum of two non-trivial Lie algebras. A classification in dimension $\leq 7$ is found in [LT22, Gon98].
On one hand, a criterion for the validity of the $\mathrm{MCP}$ on a Carnot group can be given in terms of Goh-Legendre geodesics, see Section 7.2 for details.
@@
Among the indecomposable stratified groups of dimension $\leq 7$, listed in Table 1, equipped with any choice of Carnot-Carathéodory metric and invariant measure turning it into a Carnot group:
@@ 5 $N_{5,3,1}$, $N_{5,3,2}$ $N_{5,2,3}$ $N_{5,2,1}$
Table 1. Each entry is an indecomposable stratified group of dimension $\leq 7$, see [LT22, Gon98]. Red: admits a quotient to the Martinet structure. Green: does not have non-trivial Goh-Legendre curve. Black: does not admit a quotient to Martinet and admits non-trivial Goh-Legendre curves (see Remark 1.14).
: conditional on the validity of the minimizing Sard property (see Remark 1.15). : admits non-trivial Goh curves but no Goh-Legendre ones (see Remark 7.7).
@@
We do not know whether the $\mathrm{MCP}$ is satisfied by the remaining groups in Table 1, when equipped with a Carnot structure. They do not admit a quotient to the Martinet structure, so Theorem 1.9 cannot be applied. They do have non-trivial Goh–Legendre geodesics, so we cannot use Corollary 1.12 either. Among these groups, it seems that $247H$, $247H_{1}$, $247K$, $247N$, and $247P_{1}$ admit a quotient to $N_{6,2,6}$ or $N_{6,3,1a}$. Proving that $\mathrm{MCP}$ fails for these two groups would imply the failure of the $\mathrm{MCP}$ for the others.
Remark 1.15.
@@
Remark 1.16 (Failure of the $\mathrm{MCP}$ in the pre-medium-fat case).
The notion of medium-fat distribution has been introduced in [AS99], and sub-Riemannian structures with medium-fat distribution have no non-trivial horizontal curves satisfying the Goh condition. In particular, by Corollary 1.12, any Carnot group with medium-fat distribution satisfies the $\mathrm{MCP}$. In [Rif23], the author introduces the concept of pre-medium-fat distribution, a relaxation of the medium-fat condition, allowing for non-trivial Goh-Legendre geodesics. One of the main results in [Rif23] is that sub-Riemannian structures whose distribution is pre-medium-fat satisfy the minimizing Sard property, see [Rif23, Cor. 1.3]. The author asks in [Rif23, Sec. 5.4] whether these structures do satisfy the $\mathrm{MCP}$. As an answer to that question, one can verify that the following stratified groups in Table 1 have pre-medium-fat distribution and do not satisfy any $\mathrm{MCP}$: $\mathbb{Eng}$, $N_{5,2,3}$, $N_{6,3,3}$, $N_{6,3,4}$, $357A$, $257B$, $247D$, $247E$, $247G$, and $247J$.
To justify the reduction to the indecomposable case, we end with a straightforward result for decomposable groups following from the tensorization properties of the $\mathrm{MCP}$ (see Corollary 7.8).
@@
1.8. Genericity results in high dimension and step
It is well-known that, generically, sub-Riemannian structures with constant rank $k\geq 3$ do not have non-trivial abnormal geodesics. We refer to [CJT06] for precise definitions.
Theorem 1.18 (Chitour-Jean-Trélat [CJT06]).
Let $M$ be a smooth manifold. Let $k\geq 3$ be a positive integer, and let $\mathcal{G}_{k}$ be the set of pairs $(\mathscr{D},g)$ on $M$, where $\mathscr{D}$ is a rank $k$ distribution on $M$, and $g$ is a metric tensor on $\mathscr{D}$, endowed with the Whitney $C^{\infty}$ topology. There exists an open dense subset $W_{k}$ of $\mathcal{G}_{k}$ such that, for every element of $W_{k}$, the corresponding sub-Riemannian structure does not admit non-trivial abnormal geodesics.
It is perhaps surprising that, generically, even with no non-trivial abnormal geodesics, the $\mathrm{MCP}$ fails if the dimension is large enough. The next result corresponds to Theorem 8.1.
Theorem 1.19 (Generic failure of the $\mathrm{MCP}$ for rank $3$ and high dimension).
Assume that the minimizing Sard property holds for Carnot groups. In the same setting of LABEL:{gen:genideal}, if
$\dim M\geq(k-1)\left(\frac{k^{2}}{3}+\frac{5k}{6}+1\right),$ (9)
then there exists an open dense subset $W_{k}^{\prime}\subseteq W_{k}$ of $\mathcal{G}_{k}$ such that for every element of $W_{k}^{\prime}$, the corresponding sub-Riemannian structure does not admit non-trivial abnormal geodesics and does not satisfy the $\mathrm{MCP}(K,N)$ for any $K\in\mathbb{R}$, $N\in[1,\infty)$ and any choice of smooth measure.
Remark 1.20 (Relation with generically nonsubanalytic dimensions).
The lower bound (9) is the same one that appeared in [AG01] in relation with the loss of subanaliticity of the squared distance. In fact, the underlying reason is the same: generically if (9) is true, the tangent cone at any point must admit a quotient to the Martinet structure. We also note that the lower bound (9) is not sharp and can be improved, but this is out of the scope of the current paper.
We also record the following result, corresponding to Theorem 8.2. (Note that for distributions of rank $k=2$, one does not have an analogue to Theorem 1.18; this is related with the fact that in this case the Goh condition is automatically verified by any abnormal geodesic).
Theorem 1.21 (Failure of the $\mathrm{MCP}$ for rank $2$).
Let $(M,\mathsf{d},\mathsf{m})$ be a sub-Riemannian metric measure space with constant rank $k=2$. Assume that there exists $x\in M$ such that
$\dim\mathscr{D}^{3}_{x}-\dim\mathscr{D}^{2}_{x}\geq 1.$ (10)
Assume that the minimizing Sard property holds for all Carnot groups of step $\leq s(x)$. Then $(M,\mathsf{d},\mathsf{m})$ does not satisfy the $\mathrm{MCP}(K,N)$ for any $K\in\mathbb{R}$ and $N\in[1,\infty)$.
Assumption (10) means that the iterated distribution at step $3$ at $x$ contains at least one new direction, see Appendix A for definition.
@@ 1.9. Closing thoughts
If a sub-Riemannian space does not have Goh-Legendre geodesics (including the trivial one), then, by the results of Agrachev and Lee in [AL09], the squared distance function is locally Lipschitz in charts. Consequently, thanks to the work of Badreddine and Rifford in [BR20a], and assuming that the structure is real-analytic, it must, locally, satisfy the $\mathrm{MCP}$. If Goh-Legendre geodesics are present, the $\mathrm{MCP}$ (and local Lipschitz regularity) may fail, and indeed they do fail for many examples in Table 1.
@@
Let $(M,\mathsf{d},\mathsf{m})$ be a sub-Riemannian metric measure space. If there is a strong-Goh-Legendre geodesic $\gamma$, then $\mathsf{d}^{2}:M\times M\to\mathbb{R}_{+}$ fails to be Lipschitz in charts in a neighbourhood of $\gamma$, and the $\mathrm{MCP}(K,N)$ does not hold for all $K\in\mathbb{R}$ and $N\in[1,\infty)$.
Using Corollary 1.17, we can produce Carnot groups that have non-trivial Goh-Legendre geodesics and yet satisfy the $\mathrm{MCP}$, by taking products of factors with no non-trivial Goh-Legendre geodesics. They are not strong Goh-Legendre ones though, supporting the conjecture.
@@
We remark that a trivial geodesic is Goh-Legendre if and only if the step of the structure at that point is at least $3$ – of course such geodesic cannot be a strong-Goh-Legendre one. The $\mathrm{MCP}$ is not always lost in presence of such geodesics, but local Lipschitzianity of $\mathsf{d}^{2}$ is lost in a neighbourhood of that point [AL09].
The absence of (non-trivial) Goh-Legendre geodesics also implies, for real-analytic structures, that the corresponding distance function is subanalytic, as proven by Agrachev and Gauthier in [AG01]. On the contrary, akin to what happens for the $\mathrm{MCP}$ and, of course, local Lipschitzianity, subanalyticity is lost when a suitable Martinet quotient do exist. All these properties – local Lipschitzianity, $\mathrm{MCP}$ and subanalyticity – (or their failure) seem to be a manifestation of the compactness of the set of normal geodesics of a fixed length (or lack thereof). However, a direct connection between them is currently missing.
@@ 1.10. Structure of the paper
Section 2 contains an overview of optimal transport on metric measure spaces and their quotients by isometric group actions. In Section 3, we introduce the weaker essentially non-branching condition that we require, and discuss its relation with the $\mathrm{MCP}$. We then address the stability of the $\mathrm{MCP}$ under discrete and compact group actions in Sections 4 and 5, respectively. In Section 6, we present Carnot homogeneous spaces and we prove the factorization argument. We put everything together in Section 7, proving our main theorem concerning the $\mathrm{MCP}$ of quotients of Carnot homogeneous spaces and its consequences. In Section 8, we prove the generic failure of $\mathrm{MCP}$ in high dimension and rank. In Section 9, we prove the failure of the $\mathrm{MCP}$ for the Martinet structure. We end with Section 10, where we comment on the failure of variants of the $\mathrm{MCP}$, such as its entropic version and Milman’s $\mathrm{QCD}$.